OpenStudy (anonymous):

Given the function f(x) = 3(x+2) - 4, solve for the inverse function when x = 2.

OpenStudy (jdoe0001):

\(\bf f(x)={\color{brown}{ y}} = 3({\color{blue}{ x}}+2) - 4\qquad inverse\to {\color{blue}{ x}}= 3({\color{brown}{ y}}+2) - 4\leftarrow f^{-1}\) solve for "y" notice that to find the "inverse relation" all we do is swap about the variables

OpenStudy (anonymous):

So All You Do Is Swap X And Y And Solve ? BUt It Says When The function Is 2 So I Just Plug In 2 And Solve For Y?

OpenStudy (anonymous):

Or Am I Completely Wrong ??

OpenStudy (jdoe0001):

well... after solving for "y" what did you get for the "y"?

OpenStudy (anonymous):

I Dont Understand Where I Plug In 2

OpenStudy (anonymous):

Can You Explain For A Medal ?

OpenStudy (jdoe0001):

well. how far did you get? I'd take it from there :)

OpenStudy (anonymous):

well i solved it and i got 0 for y

OpenStudy (jdoe0001):

well o...ahemm 0..? one sec \(\bf \bf f(x)={\color{brown}{ y}} = 3({\color{blue}{ x}}+2) - 4\qquad inverse\to {\color{blue}{ x}}= 3({\color{brown}{ y}}+2) - 4\leftarrow f^{-1} \\ \quad \\ \to x+4=3(y+2)\implies \cfrac{x+4}{3}=y+2\implies \cfrac{x+4}{3}-2=y\iff f^{-1} \\ \quad \\ \textit{now let's set }x=2\qquad FOR \quad f^{-1} \\ \quad \\ f^{-1}({\color{brown}{ x}})=\cfrac{{\color{brown}{ x}}+4}{3}-2\implies f^{-1}({\color{brown}{ 2}})=\cfrac{{\color{brown}{ 2}}+4}{3}-2\)

OpenStudy (anonymous):

−4 0 4 8 theses are the choices ,.....

OpenStudy (jdoe0001):

so... what would that give you?

OpenStudy (anonymous):

2 + 4 =6 Then / BY 3 = 2 Then - 2 Which = 0

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Can You Help Me With Another ??